### v1

# Non-convergence of the Navier-Stokes equations toward the Euler equations in weak Besov spaces

Abstract
In this paper, we consider the inviscid limit problem to the higher dimensional incompressible Navier-Stokes equations in the whole space. It was proved in [16, J. Funct. Anal., 276 (2019)] that given initial data $u_{0}\in B^{s}_{p,r}$ with $1\leq r<\infty$, the solutions of the Navier-Stokes equations converge strongly in $B^{s}_{p,r}$ to the Euler equations as the viscosity parameter tends to zero. In the case when $r=\infty$, we prove the failure of the $B^{s}_{p,\infty}$-convergence of the Navier-Stokes equations toward the Euler equations in the inviscid limit.
Keywords: Navier-Stokes equations, Euler equations, Inviscid limit, Besov spaces.
MSC (2010): 35Q30, 76B03.
1 Introduction
The motion of an incompressible viscous fluid with constant density is governed by the following Navier-Stokes equations:
$\begin{cases}\partial_{t}u+u\cdot\nabla u-\varepsilon\Delta u+\nabla P=0,\\ \mathrm{div\,}u=0,\\ u(0,x)=u_{0}(x),\end{cases}$ (NS)
where $\varepsilon>0$ denotes the viscosity of the fluid, the vector field $u(t,x):[0,\infty)\times{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ stands for the velocity of the fluid, the quantity $P(t,x):[0,\infty)\times{\mathbb{R}}^{d}\to{\mathbb{R}}$ denotes the scalar pressure, and $\mathrm{div\,}u=0$ means that the fluid is incompressible.
When the viscocity vanishes ($\varepsilon=0$), then the Navier-Stokes equations (NS) reduces to the Euler equations for ideal incompressible fluid
$\begin{cases}\partial_{t}u+u\cdot\nabla u+\nabla P=0,\\ \mathrm{div\,}u=0,\\ u(0,x)=u_{0}(x).\end{cases}$ (E)
The mathematical study of both the Navier-Stokes equations (NS) and Euler equations (E) has a long and distinguished history (see Constantin’s survey [7] for more details). We do not detail the literature since it is huge and refer the readers to see the monographs of Majda-Bertozzi [22] and Bahouri-Chemin-Danchin [1] for the well-posedness results of both Navier-Stokes and Euler equations in Sobolev and Besov spaces respectively.
A classical problem in fluid mechanics is the approximation in the limit $\varepsilon\to 0$ of vanishing viscosity (also called inviscid limit or zero-viscosity limit) of solutions of the Euler equations by solutions of the incompressible Navier-Stokes equations, which is naturally associated with the physical phenomena of turbulence and of boundary layers. The vanishing viscosity problem is closely related to uniqueness of solutions to the Euler equations, because the methods used to prove uniqueness can often be applied to show vanishing viscosity. In the presence of boundaries, the zero-viscosity limit problem is more complicated and essentially difficult as it is closely related to the validity of the Prandtl equation for the formation of boundary layers. We refer to see the recent results by Maekawa [20], Constantin, Kukavica and Vicol [10, 11], Gérard et al. [15] and the references therein. For initial data that is analytic only close to the boundary of the domain, Kukavica, Vicol and Wang [26, 27] proved that the solution of the Navier-Stokes equations converges in the vanishing viscosity limit to the solution of the Euler equations.
Most of the progress has been made in the two-dimensional case without boundary (assuming the domain to be the whole plane $\mathbb{R}^{2}$ of the torus $\mathbb{T}^{2}$) since both vortex stretching and boundary effects are absent. In particular, the problem of the convergence of smooth viscous solutions of the Navier-Stokes equations to the Eulerian one as $\varepsilon\to 0$ is well understood and has been studied in many literatures, see for example [9, 29, 19]. Majda [21] showed that under the assumption $u_{0}\in H^{s}$ with $s>\frac{d}{2}+2$, the solutions to (NS) converge in $L^{2}$-norm as $\varepsilon\to 0$ to the unique solution to (E) and the convergence rate is of order $(\varepsilon t)^{\frac{1}{2}}$. Masmoudi [23] improved the result and proved the convergence in $H^{s}$-norm under the weaker assumption $u_{0}\in H^{s}$ with $s>\frac{d}{2}+1$. Next we shall recall the consideration of Yudovich solutions. Yudovich [31] established the uniqueness of solutions to the Euler equations in the space $C\left(\mathbb{R}^{+};L^{2}\left(\mathbb{R}^{2}\right)\right)\times L_{\text{loc }}^{\infty}\left(\mathbb{R}^{+};L^{2}\left(\mathbb{R}^{2}\right)\right)$ assuming that initial data lies to $L^{2}\left(\mathbb{R}^{2}\right)$ and initial vorticity belongs to $L^{p}\left(\mathbb{R}^{2}\right)\cap L^{\infty}\left(\mathbb{R}^{2}\right)$ for some $p<\infty$. For this uniqueness class, Chemin [5] showed that solutions of the Navier-Stokes equations with initial vorticity lies in $L^{2}(\mathbb{R}^{2})\cap L^{\infty}(\mathbb{R}^{2})$ converge in the zero-viscosity limit to a solution of the Euler equations. Kelliher [25] resolved the same inviscid limit as Chemin [5] with the assumption that the vorticity is unbounded. We also mention that, in the case of that initial velocity has finite energy and initial vorticity is bounded in the plane, Cozzi [12, 13] showed that the unique solutions of the Navier-Stokes equations converge to the unique solution of the Euler equations in the $L^{\infty}$-norm uniformly over finite time as viscosity approaches zero. In the case of two dimensional (torus or the whole space) and rough initial data, by taking greater advantage of vorticity formulation, more beautiful results were obtained quantitatively (see for example [2, 14, 6, 8, 25, 32] and the references therein).
Next, we review some results on the inviscid limit under the framework of Besov spaces. In dimension two, Hmidi and Kerrani [17] proved that (NS) is globally well-posed in Besov space $B^{2}_{2,1}$, with uniform bounds on the viscosity and obtained that the convergence rate of the inviscid limit is of order $\varepsilon t$ for vanishing viscosity. Subsequently, in [18], they further generalized to other Besov spaces $B^{2/p+1}_{p,1}$ with convergence in $L^{p}$. For more details, we also refer to see [24, Section 3.4]. In dimension three, a similar result was obtained in [30] for axis-symmetric flows without swirl. Bourgain and Li [3, 4] employed a combination of Lagrangian and Eulerian techniques to obtain strong local ill-posedness results in borderline Besov spaces $B^{d/p+1}_{p,r}$ for $1\leq p<\infty$ and $1<r\leq\infty$ when $d=2,3$. Guo-Li-Yin [16] solved the inviscid limit in the same topology. Precisely speaking, they obtained
Theorem 1.1 ([16]).
Let $d\geq 2$ and $\varepsilon\in[0,1]$. Assume that $(s,p,r)$ satisfies
$s>\frac{d}{p}+1,\;(p,r)\in[1,\infty]\times[1,\infty)\quad\mbox{ or }$
$s=\frac{d}{p}+1,\;(p,r)\in[1,\infty]\times\{1\}.$
Given initial data $u_{0}\in B^{s}_{p,r}(\mathbb{R}^{d})$, there exists $T=T(s,p,r,d)>0$ such that (NS) has a unique solution $u^{\rm NS}_{\varepsilon}(t,u_{0})\in C([0,T];B^{s}_{p,r})$. Moreover, there holds
(1)  Uniform bounds: there exists $C=C(s,p,r,d)>0$ such that
$\left\|u^{\rm NS}_{\varepsilon}(t,u_{0})\right\|_{L_{T}^{\infty}B^{s}_{p,r}}\leq C,\quad\forall\ \varepsilon\in[0,1].$
Moreover, if $u_{0}\in B^{\gamma}_{p,r}$ with $\gamma>s$, there exists $C_{2}=C_{2}(\gamma,s,p,r,d)>0$ such that
$\left\|u^{\rm NS}_{\varepsilon}(t,u_{0})\right\|_{L_{T}^{\infty}B^{\gamma}_{p,r}}\leq C_{2}\|u_{0}\|_{B^{\gamma}_{p,r}}.$
(2) Inviscid limit: $\lim\limits_{\varepsilon\downarrow 0}\left\|u^{\rm NS}_{\varepsilon}(t,u_{0})-u^{\rm E}(t,u_{0})\right\|_{L_{T}^{\infty}B^{s}_{p,r}}=0,$ namely, for any $u_{0}\in B^{s}_{p,r}$, $\forall\eta>0$, $\exists\ \varepsilon_{0}=\varepsilon_{0}(u_{0},\eta,T)$, $\forall\varepsilon\leq\varepsilon_{0}$, such that
$\left\|u^{\rm NS}_{\varepsilon}(t,u_{0})-u^{\rm E}(t,u_{0})\right\|_{L_{T}^{\infty}B^{s}_{p,r}}\leq\eta.$
However, we should be noted that it left an open problem for the end-point case $r=\infty$ in the inviscid limit of the Navie-Stokes equations in Besov spaces. Precisely speaking, the statement (2) is a consequence of the strong convergence of the solutions to (NS) with the given initial data belonging in $B^{s}_{p,r}$ with $1\leq r<\infty$. An interesting Question appears:
For any $u_{0}\in B^{s}_{p,\infty}$, in the inviscid limit $\varepsilon\to 0$, whether or not the $B^{s}_{p,\infty}$-convergence
$u^{\rm NS}_{\varepsilon}(t,u_{0})\rightarrow u^{\rm E}(t,u_{0})\quad\text{strongly in}\quad L_{T}^{\infty}B^{s}_{p,\infty}$
can be established?
In this paper, we will give a negative answer: there can be no strong convergence in $B^{s}_{p,\infty}$, i.e.,
$u^{\rm NS}_{\varepsilon}(t,u_{0})\nrightarrow u^{\rm E}(t,u_{0})\quad\text{strongly in}\quad L_{T}^{\infty}B^{s}_{p,\infty}.$
For any $R>0$, from now on, we denote any bounded subset $U_{R}$ in $B^{s}_{p,\infty}(\mathbb{R}^{d})$ by
$U_{R}:=\left\{\phi\in B^{s}_{p,\infty}(\mathbb{R}^{d}):\|\phi\|_{B^{s}_{p,\infty}(\mathbb{R}^{d})}\leq R,\;\mathrm{div\,}\phi=0\right\}$
and also
[…]
Now we state main result.
Theorem 1.2.
Let $d\geq 2$ and $\varepsilon\in[0,1]$. Assume that $(s,p)$ satisfies
$s>\frac{d}{p}+1\quad\text{with}\quad p\in[1,\infty].$ (1.1)
For any $u_{0}\in U_{R}\subseteq B^{s}_{p,\infty}(\mathbb{R}^{d})$, then a family of solutions $\left\{u^{\rm NS}_{\varepsilon}(t,u_{0})\right\}_{\varepsilon>0}$ to (NS) do not converge to the solution $u^{\rm E}(t,u_{0})$ of (E) in $B^{s}_{p,\infty}(\mathbb{R}^{d})$. More precisely, there exists a sequence initial data $u_{0}\in U_{R}$ such that for a short time $T_{0}\leq T$
$\liminf_{\varepsilon\downarrow 0}\left\|u^{\rm NS}_{\varepsilon}(t,u_{0})-u^{\rm E}(t,u_{0})\right\|_{L^{\infty}_{T_{0}}B^{s}_{p,\infty}}\geq\eta_{0},$
with some positive constant $\eta_{0}$ depending on $d$, but independent of $\varepsilon$.
Remark 1.1.
To the best of our knowledge, Theorem 1.2 is new. In fact, it was proved in [16] that given initial data $u_{0}\in B^{s}_{p,r}$ with $1\leq r<\infty$ and for some $T>0$, the solutions of the Navier-Stokes equations converge strongly in $L^{\infty}_{T}B^{s}_{p,r}$ to the Euler equations as the viscosity parameter tends to zero while our Theorem 1.2 shows the non-convergence of the solutions for the Navier-Stokes equations toward the Euler equations in the inviscid limit under the framework of weak Besov spaces $B^{s}_{p,\infty}$.
Strategies to the proof of Theorem 1.2.
Due to the incompressible condition, the pressure can be eliminated from (NS) and (E). In fact, applying the Leray operator $\mathcal{P}$ to (NS) and (E), respectively, then we have
$\begin{cases}\partial_{t}u-\varepsilon\Delta u+\mathcal{P}(u\cdot\nabla u)=0,\\ u(0,x)=u_{0}(x),\end{cases}$
and
$\begin{cases}\partial_{t}u+\mathcal{P}(u\cdot\nabla u)=0,\\ u(0,x)=u_{0}(x).\end{cases}$
From Duhamel’s principle, it follows that
$u^{\rm NS}_{\varepsilon}(t,u_{0})=e^{t\varepsilon\Delta}u_{0}-\int_{0}^{t}e^{(t-\tau)\varepsilon\Delta}\mathcal{P}(u^{\rm NS}_{\varepsilon}\cdot\nabla u^{\rm NS}_{\varepsilon})\mathrm{d}\tau,$ (1.2)
$u^{\rm E}(t,u_{0})=u_{0}-\int_{0}^{t}\mathcal{P}\left(u^{\rm E}\cdot\nabla u^{\rm E}\right)\mathrm{d}\tau.$ (1.3)
Step 1: At a first sight, to ensure that the difference between the solutions $u^{\rm NS}_{\varepsilon}$ to (NS) and $u^{\rm E}$ to (E) in the $B^{s}_{p,\infty}$-topology is bounded below by a positive constant at any later time, we need to construct the initial data sequence $u_{0}\in B^{s}_{p,\infty}$ to satisfy that the difference between the solutions $e^{t\varepsilon\Delta}u_{0}$ and $u_{0}$ in the $B^{s}_{p,\infty}$-topology is bounded below by a positive constant at any later time. In order to see this, let us introduce the following toy-model:
A toy-model: 1-D heat equation. Let $\varepsilon_{n}:=2^{-2n}$ with $n\gg 1$. The Cauchy problem
$\begin{cases}\partial_{t}u-\varepsilon_{n}\partial^{2}_{xx}u=0,\\ u(0,x)=u_{0}(x)\in B^{s}_{p,\infty}(\mathbb{R}),\end{cases}$
has a unique explicit solution $u^{\varepsilon_{n}}(x,t)=e^{\varepsilon_{n}t\partial^{2}_{xx}}u_{0}(x)$, where the heat kernel operator is given by $e^{\varepsilon_{n}t\partial^{2}_{xx}}:=\mathcal{F}^{-1}e^{-\varepsilon_{n}t|\xi|^{2}}\mathcal{F}.$
Let $\phi$ be given in subsection 3.1 and set $u_{0}(x)=2^{-ns}\phi(x)\cos\left(\frac{17}{12}2^{n}x\right)$, then $u^{\varepsilon_{n}}\nrightarrow u_{0}$ in $B^{s}_{p,\infty}(\mathbb{R})$. In fact, for $n\gg 1$, one has for $t>0$
$\left\|e^{\varepsilon_{n}t\partial^{2}_{xx}}u_{0}-u_{0}\right\|_{B^{s}_{p,\infty}}\geq\left(e^{9t/16}-1\right)\left\|\phi(x)\cos\left(\frac{17}{12}2^{n}x\right)\right\|_{L^{p}}\nrightarrow 0\quad\text{as}\;n\to\infty.$
Step 2: It remains to prove that the error term
$\int_{0}^{t}e^{(t-\tau)\varepsilon\Delta}\mathcal{P}(u^{\rm NS}_{\varepsilon}\cdot\nabla u^{\rm NS}_{\varepsilon})\mathrm{d}\tau-\int_{0}^{t}\mathcal{P}\left(u^{\rm E}\cdot\nabla u^{\rm E}\right)\mathrm{d}\tau$
is smaller than the leading term $e^{t\varepsilon\Delta}u_{0}-u_{0}$. We can decompose the above error term into four terms (see (3.4) in subsection 3.4) and estimate them by subtle argument (see Propositions 3.3-3.4 in subsection 3.3).
Organization of our paper.
In Section 2, we list some notations and known results which will be used in the sequel. We prove Theorem 1.2 in Section 3 by dividing it into several parts:
(1) Construction and Estimation of initial data (see subsection 3.1);
(2) Estimation of leading term (see subsection 3.2);
(3) Estimation of error terms (see subsection 3.3);
(4) Non-convergence (see subsection 3.4).
2 Preliminaries
We will use the following notations throughout this paper.
  • For $X$ a Banach space and $I\subset\mathbb{R}$, we denote by $\mathcal{C}(I;X)$ the set of continuous functions on $I$ with values in $X$. Sometimes we will denote $L^{p}(0,T;X)$ by $L_{T}^{p}X$.
  • The symbol $\mathrm{A}\approx\mathrm{B}$ means that there is a uniform positive “harmless” constant $C$ independent of $\mathrm{A}$ and $\mathrm{B}$ such that $C^{-1}\mathrm{B}\leq\mathrm{A}\leq C\mathrm{B}$.
  • Let us recall that for all $u\in\mathcal{S}^{\prime}$, the Fourier transform $\mathcal{F}u$, also denoted by $\widehat{u}$, is defined by
    $\mathcal{F}u(\xi)=\widehat{u}(\xi)=\int_{\mathbb{R}^{d}}e^{-\mathrm{i}x\cdot\xi}u(x)\mathrm{d}x\quad\text{for any}\;\xi\in\mathbb{R}^{d}.$
  • The inverse Fourier transform allows us to recover $u$ from $\widehat{u}$:
    $u(x)=\mathcal{F}^{-1}\widehat{u}(x)=(2\pi)^{-d}\int_{\mathbb{R}^{d}}e^{\mathrm{i}x\cdot\xi}\widehat{u}(\xi)\mathrm{d}\xi.$
  • We denote the projection
    $\mathcal{P}:L^{p}(\mathbb{R}^{d})\rightarrow L_{\sigma}^{p}(\mathbb{R}^{d})\equiv\overline{\left\{f\in\mathcal{C}^{\infty}_{0}(\mathbb{R}^{d});{\rm{div}}f=0\right\}}^{\|\cdot\|_{L^{p}(\mathbb{R}^{d})}},\quad p\in(1,\infty),$
    $\mathcal{Q}=\mathrm{Id}-\mathcal{P}.$
    In $\mathbb{R}^{d}$, $\mathcal{P}$ can be defined by $\mathcal{P}=\mathrm{Id}+(-\Delta)^{-1}\nabla{\rm{div}}$, or equivalently, $\mathcal{P}=(\mathcal{P}_{ij})_{1\leqslant i,j\leqslant d}$, where $\mathcal{P}_{ij}\equiv\delta_{i,j}+R_{i}R_{j}$ with $\delta_{i,j}$ being the Kronecker delta ($\delta_{i,j}=0$ for $i\neq j$ and $\delta_{i,i}=0$) and $R_{i}$ being the Riesz transform with symbol $-\mathrm{i}\xi_{1}/|\xi|$. Obviously, $\mathcal{Q}=-(-\Delta)^{-1}\nabla{\rm{div}}$, and if $\mathord{{\rm div}}\,u=\mathord{{\rm div}}\,v=0$, it holds that $\mathcal{Q}(u\cdot\nabla v)=\mathcal{Q}(v\cdot\nabla u).$
Next, we will recall some facts about the Littlewood-Paley decomposition and the nonhomogeneous Besov spaces (see [1] for more details). Choose a radial, non-negative, smooth function $\vartheta:\mathbb{R}^{d}\mapsto[0,1]$ such that
  • ${\rm{supp}}\;\vartheta\subset B(0,4/3)$;
  • $\vartheta(\xi)\equiv 1$ for $|\xi|\leq 3/4$.
Setting $\varphi(\xi):=\vartheta(\xi/2)-\vartheta(\xi)$, then we deduce that $\varphi$ has the following properties
  • ${\rm{supp}}\;\varphi\subset\left\{\xi\in\mathbb{R}^{d}:3/4\leq|\xi|\leq 8/3\right\}$;
  • $\varphi(\xi)\equiv 1$ for $24/3\leq|\xi|\leq 3/$ 2;
  • $\vartheta(\xi)+\sum_{j\geq 0}\varphi(2^{-j}\xi)=1$ for any $\xi\in\mathbb{R}^{d}$;
  • $\sum_{j\in\mathbb{Z}}\varphi(2^{-j}\xi)=1$ for any $\xi\in\mathbb{R}^{d}\setminus\{0\}$.
For any $u\in\mathcal{S^{\prime}}(\mathbb{R}^{d})$, the nonhomogeneous and homogeneous dyadic blocks are defined as follows
$\Delta_{j}u=0,\;\text{if}\;j\leq-2;$
$\Delta_{-1}u=\vartheta(D)u;$
$\Delta_{j}u=\varphi(2^{-j}D)u,\;\;\text{if}\;j\geq 0,$
and for any $u\in\mathcal{S}^{\prime}_{h}(\mathbb{R}^{d})$
$\Delta_{j}u=\varphi(2^{-j}D)u,\;\;\text{if}\;j\in\mathbb{Z},$
where the pseudo-differential operator is defined by $\sigma(D):u\to\mathcal{F}^{-1}(\sigma\mathcal{F}u)$ and $\mathcal{S}^{\prime}_{h}$ is given by
$\mathcal{S}^{\prime}_{h}:=\Big{\{}u\in\mathcal{S^{\prime}}(\mathbb{R}^{d}):\;\lim_{j\rightarrow-\infty}\|\vartheta(2^{-j}D)u\|_{L^{\infty}}=0\Big{\}}.$
Remark 2.1.
Due to the fact $\varphi(\xi)\equiv 1$ for $\frac{4}{3}\leq|\xi|\leq\frac{3}{2}$, for $-1\leq j,n\in\mathbb{Z}$, we have
${\Delta_{n}(f_{j})=\mathcal{F}^{-1}\left(\varphi(2^{-n}\cdot)\widehat{f}_{j}\right)=}\begin{cases}f_{j},&\text{if}\;n=j,\\ 0,&\text{if}\;n\neq j.\end{cases}$
We recall the definition of the Besov Spaces and norms.
Definition 2.1 (Besov Spaces, [1]).
Let $s\in\mathbb{R}$ and $(p,r)\in[1,\infty]^{2}$. The nonhomogeneous Besov spaces are defined as follows
$B_{p,r}^{s}(\mathbb{R}^{d})=\left\{f\in\mathcal{S}^{\prime}(\mathbb{R}^{d}):\;\|f\|_{B_{p,r}^{s}(\mathbb{R}^{d})}<\infty\right\},\quad\text{where}$
$\|f\|_{B^{s}_{p,r}(\mathbb{R}^{d})}:=$ $\left(\sum_{j\geq-1}2^{sjr}\|\Delta_{j}f\|^{r}_{L^{p}(\mathbb{R}^{d})}\right)^{1/r},$ if $1\leq r<\infty$,
$\|f\|_{B^{s}_{p,r}(\mathbb{R}^{d})}:=$ $\sup_{j\geq-1}\left(2^{sj}\|\Delta_{j}f\|_{L^{p}(\mathbb{R}^{d})}\right),$ if $r=\infty$.
The homogeneous Besov spaces are defined as follows
$\dot{B}_{p,r}^{s}(\mathbb{R}^{d})=\Big{\{}f\in\mathcal{S}^{\prime}_{h}(\mathbb{R}^{d}):\;\|f\|_{\dot{B}_{p,r}^{s}(\mathbb{R}^{d})}<\infty\Big{\}},\quad\text{where}$
$\|f\|_{\dot{B}^{s}_{p,r}(\mathbb{R}^{d})}:=$ $\left(\sum_{j\in\mathbb{Z}}2^{sjr}\|\dot{\Delta}_{j}f\|^{r}_{L^{p}(\mathbb{R}^{d})}\right)^{1/r},$ if $1\leq r<\infty$,
$\|f\|_{\dot{B}^{s}_{p,r}(\mathbb{R}^{d})}:=$ $\sup_{j\in\mathbb{Z}}\left(2^{sj}\|\dot{\Delta}_{j}f\|_{L^{p}(\mathbb{R}^{d})}\right),$ if $r=\infty$.
We remark that, for any $s>0$ and $(p,r)\in[1,\infty]^{2}$, then $B^{s}_{p,r}(\mathbb{R}^{d})=\dot{B}^{s}_{p,r}(\mathbb{R}^{d})\cap L^{p}(\mathbb{R}^{d})$ and
$\|f\|_{B^{s}_{p,r}}\approx\|f\|_{L^{p}}+\|f\|_{\dot{B}^{s}_{p,r}}.$
The following Bernstein’s inequalities will be used in the sequel.
Lemma 2.1 (Bernstein’s inequalities, [1]).
Let $\mathcal{B}$ be a ball and $\mathcal{C}$ be an annulus. There exists a positive constant $C$ such that for all $k\in\mathbb{N}\cup\{0\}$, any $\lambda\in\mathbb{R}^{+}$ and any function $f\in L^{p}$ with $1\leq p\leq q\leq\infty$
${\rm{supp}}\ \widehat{f}\subset\lambda\mathcal{B}\;\Rightarrow\;\|D^{k}f\|_{L^{q}}\leq C^{k+1}\lambda^{k+(\frac{d}{p}-\frac{d}{q})}\|f\|_{L^{p}},$
${\rm{supp}}\ \widehat{f}\subset\lambda\mathcal{C}\;\Rightarrow\;C^{-k-1}\lambda^{k}\|f\|_{L^{p}}\leq\|D^{k}f\|_{L^{p}}\leq C^{k+1}\lambda^{k}\|f\|_{L^{p}}.$
Finally we recall the following product law which will be used often in the sequel.
Lemma 2.2 (Product estimation, [1]).
Assume $(s,p)$ satisfies (1.1). $B^{s}_{p,\infty}(\mathbb{R}^{d})$ is a Banach algebra. For any $r\in[1,\infty]$, then there exists a constant $C>0$ such that
$\|fg\|_{B^{s}_{p,r}}\leq C\left(\|f\|_{L^{\infty}}\|g\|_{B^{s}_{p,r}}+\|g\|_{L^{\infty}}\|f\|_{B^{s}_{p,r}}\right),\quad\forall f,g\in L^{\infty}\cap B^{s}_{p,r}.$
Furthermore, due to the embedding $B^{s-1}_{p,r}(\mathbb{R}^{d})\hookrightarrow L^{\infty}(\mathbb{R}^{d})$, there holds
$\|f\cdot\nabla g\|_{B^{s-1}_{p,r}}\leq C\|f\|_{B^{s-1}_{p,r}}\|g\|_{B^{s}_{p,r}},\quad\forall(f,g)\in B^{s-1}_{p,r}\times B^{s}_{p,r}.$
3 Proof of Theorem 1.2
In this section, we prove Theorem 1.2 by dividing it into several parts: (1) Construction and Estimation of initial data; (2) Estimation of leading term; (3) Estimation of errors terms; (4) Non-convergence.
Before proving Theorem 1.2, we recall the well-posedness result. Assume $(s,p)$ satisfies the condition (1.1). It is well-known, by the classical well-posedness theory, that for fixed $\varepsilon>0$, there exists $T=T(\|u_{0}\|_{B^{s}_{p,\infty}})>0$ independent of $\varepsilon$ such that (NS) has a unique solution $u$ in $L^{\infty}([0,T];B^{s}_{p,\infty})$.
3.1 Construction of Initial Data
We need to introduce smooth, radial cut-off functions to localize the frequency region. Let $\widehat{\phi}\in\mathcal{C}^{\infty}_{0}(\mathbb{R})$ be an even, real-valued and non-negative function on $\mathbb{R}$ and satisfy
${\widehat{\phi}(\xi)=}\begin{cases}1,&\mathrm{if}\;|\xi|\leq\frac{1}{4^{d}},\\ 0,&\mathrm{if}\;|\xi|\geq\frac{1}{2^{d}}.\end{cases}$
Remark 3.1.
For any $p\in[1,\infty]$, there exists two positive constants $c_{1}$ and $C_{1}$, we have
$c_{1}\leq\|\phi\|_{L^{p}(\mathbb{R})}\leq C_{1}.$
Lemma 3.1.
Let $s\in\mathbb{R}$ and $p\in[1,\infty]$. For some fixed $\delta\in(0,1)$ which will be determined later, we define the divergence-free vector field $u_{0}$ by
$u_{0}(x):=\delta\left(-\partial_{2},\;\partial_{1},\;0,\;\ldots,\;0\right)\sum_{j=3}^{\infty}2^{-j(s+1)}f_{j}(x),\quad\text{where}$
$f_{j}(x):=\phi(x_{1})\cos\left(\frac{17}{12}2^{j}x_{1}\right)\prod_{i=2}^{d}\phi(x_{i}).$
Then there exists some sufficiently large $n\in\mathbb{Z}^{+}$ such that
$\|u_{0}\|_{B^{s}_{p,\infty}}\thickapprox\delta\quad\text{and}\quad 2^{ns}\|\Delta_{n}u_{0}\|_{L^{p}}\thickapprox\delta.$ (3.4)
Proof.
We assume that $p\in[1,\infty)$ without loss of generality. By easy computations, we deduce that
$\widehat{f_{j}}(\xi)=\frac{1}{2}\left[\widehat{\phi}\left(\xi_{1}+\frac{17}{12}2^{j}\right)+\widehat{\phi}\left(\xi_{1}-\frac{17}{12}2^{j}\right)\right]\prod_{i=2}^{d}\widehat{\phi}\left(\xi_{i}\right),$
which implies that
$\mathrm{supp}\ \widehat{f_{j}}$$\subset\left\{\xi\in\mathbb{R}^{d}:\ \frac{33}{24}2^{j}\leq|\xi|\leq\frac{35}{24}2^{j}\right\}.$
Noticing that $\Delta_{k}(f_{j})=\delta_{k,j}f_{j}$ for some $k\geq-1$ (see Remark 2.1), we deduce that
[…]
which implies that
$\|u_{0}\|_{{B}_{p,\infty}^{s}(\mathbb{R}^{d})}$ $=\sup_{k\geq-1}\left(2^{ks}\|\Delta_{k}u_{0}\|_{L^{p}(\mathbb{R}^{d})}\right)$
$=\delta\sup_{k\geq 0}\left(2^{-k}\|\partial_{1}f_{k}\|_{L^{p}(\mathbb{R}^{d})}+2^{-k}\|\partial_{2}f_{k}\|_{L^{p}(\mathbb{R}^{d})}\right)\thickapprox\delta,$
and
$2^{ns}\|\Delta_{n}u_{0}\|_{L^{p}(\mathbb{R}^{d})}$ $=\delta 2^{-n}\left(\|\partial_{1}f_{n}\|_{L^{p}(\mathbb{R}^{d})}+\|\partial_{2}f_{n}\|_{L^{p}(\mathbb{R}^{d})}\right)\thickapprox\delta,$
where we have used Remark 3.1. We complete the proof of Lemma 3.1. ∎
3.2 Estimation of Leading Term
Now we establish the following proposition involving the leading term $e^{t\varepsilon\Delta}u_{0}-u_{0}$ which will affect the non-convergence of solutions to (NS) in the proof of Theorem 1.2.
From now on, we set $\varepsilon:=2^{-n}$ with $n\gg 1$.
Proposition 3.1.
Let $u_{0}$ be given by Lemma 3.1. Assume that $(s,p)$ satisfies (1.1). For $t\in(0,1)$, we have for some positive constants $c_{1},C_{1},C_{2}$ independent of $t,\varepsilon$ and $\delta$
$\delta t2^{n}\left(c_{1}-C_{1}t2^{n}\right)\leq 2^{ns}\left\|\Delta_{n}\big{(}e^{t\varepsilon\Delta}u_{0}-u_{0}\big{)}\right\|_{L^{p}}\leq C_{2}\delta 2^{n}t.$
Proof.
By the Newton-Leibniz formula, one has
$e^{t\varepsilon\Delta}u_{0}-u_{0}=\int_{0}^{t}\left(e^{\tau\varepsilon\Delta}\varepsilon\Delta u_{0}\right)\mathrm{d}\tau.$ (3.5)
From (3.5), then by Bernstein’s inequality (see Lemma 2.1), we obtain from (3.4) that for large $n$
$2^{ns}\big{\|}\Delta_{n}\big{(}e^{t\varepsilon\Delta}u_{0}-u_{0}\big{)}\big{\|}_{L^{p}}$ $\leq t2^{n}2^{ns}\big{\|}{\Delta}_{n}\left(\varepsilon\Delta u_{0}\right)\big{\|}_{L^{p}}$
$\thickapprox t2^{n}2^{ns}\big{\|}{\Delta}_{n}u_{0}\big{\|}_{L^{p}}$
$\thickapprox t\delta 2^{n}.$
Combining the above, we obtain that for large $n$
$2^{ns}\big{\|}\Delta_{n}\big{(}e^{t\varepsilon\Delta}u_{0}-u_{0}\big{)}\big{\|}_{L^{p}}\geq$ $~{}t2^{ns}\big{\|}{\Delta}_{n}\varepsilon\Delta u_{0}\big{\|}_{L^{p}}-2^{ns}\int_{0}^{t}\left\|{\Delta}_{n}\varepsilon\Delta\big{(}e^{\tau\varepsilon\Delta}u_{0}-u_{0}\big{)}\right\|_{L^{p}}\mathrm{d}\tau$
$\geq$ $~{}ct2^{ns}2^{n}\big{\|}{\Delta}_{n}u_{0}\big{\|}_{L^{p}}-C2^{n}2^{ns}\int_{0}^{t}\left\|{\Delta}_{n}\big{(}e^{\tau\varepsilon\Delta}u_{0}-u_{0}\big{)}\right\|_{L^{p}}\mathrm{d}\tau$
$\geq$ $~{}c_{1}t\delta 2^{n}-C_{1}t^{2}\delta 2^{2n}$
$\geq$ $~{}\delta t2^{n}\left(c_{1}-C_{1}t2^{n}\right).$
Thus we complete the proof of Proposition 3.1. ∎
3.3 Estimation of Error Terms
The following propositions involving the perturbation of solutions will play a crucial role in the proof of Theorem 1.2.
Proposition 3.2.
Let $u_{0}$ be given by Lemma 3.1. Assume that $(s,p)$ satisfies (1.1). For $t\in(0,1)$, we have for some positive constant $C$ independent of $t,\varepsilon$ and $\delta$
$\left\|e^{t\varepsilon\Delta}u_{0}\cdot\nabla e^{t\varepsilon\Delta}u_{0}-u_{0}\cdot\nabla u_{0}\right\|_{\dot{B}^{s-1}_{p,\infty}}\leq C\delta^{2},$
$\left\|\left(e^{t\varepsilon\Delta}-\mathrm{Id}\right)\left(u_{0}\cdot\nabla u_{0}\right)\right\|_{\dot{B}^{s-1}_{p,\infty}}\leq C\delta^{2}.$
Proof.
Noticing that $\mathrm{div\,}u_{0}=0$ and using Lemmas 2.2-3.1, we have
$\left\|e^{t\varepsilon\Delta}u_{0}\cdot\nabla e^{t\varepsilon\Delta}u_{0}-u_{0}\cdot\nabla u_{0}\right\|_{\dot{B}^{s-1}_{p,\infty}}\leq$ $~{}C\left\|e^{t\varepsilon\Delta}u_{0}\otimes e^{t\varepsilon\Delta}u_{0}-u_{0}\otimes u_{0}\right\|_{B^{s}_{p,\infty}}$
$\leq$ $~{}C\left\|u_{0}\right\|^{2}_{B^{s}_{p,\infty}}$
$\leq$ $~{}C\delta^{2},$
and
$\left\|\left(e^{t\varepsilon\Delta}-\mathrm{Id}\right)\left(u_{0}\cdot\nabla u_{0}\right)\right\|_{\dot{B}^{s-1}_{p,\infty}}\leq$ $~{}C\left\|u_{0}\otimes u_{0}\right\|_{B^{s}_{p,\infty}}\leq C\delta^{2}.$
This finishes the proof of Proposition 3.2. ∎
Proposition 3.3.
Let $u_{0}$ be given by Lemma 3.1. Assume that $(s,p)$ satisfies (1.1). For $t\in(0,1)$, we have for some positive constant $C$ independent of $t,\varepsilon$ and $\delta$
$\left\|u^{\rm E}(t,u_{0})-u_{0}+t\mathcal{P}\left(u_{0}\cdot\nabla u_{0}\right)\right\|_{\dot{B}^{s-2}_{p,\infty}}\leq Ct^{2}\delta^{3}.$
Proof.
From now on, we set $u^{\rm E}_{t}=u^{\rm E}(t,u_{0})$ for simplicity. Due to the fact $B^{s}_{p,\infty}\hookrightarrow\rm Lip$, we know that there exists a positive time $T=T(\|u_{0}\|_{B^{s}_{p,\infty}})$ such that
$\|u^{\rm E}_{t}\|_{L^{\infty}_{T}B^{s}_{p,\infty}}\leq C\|u_{0}\|_{B^{s}_{p,\infty}}\leq C\delta.$
By the Mean value theorem, from (E), one has
$u^{\rm E}_{t}-u_{0}=-\int_{0}^{t}\mathcal{P}\left(u^{\rm E}_{\tau}\cdot\nabla u^{\rm E}_{\tau}\right)\mathrm{d}\tau.$ (3.6)
By Minkowski’s inequality and Lemma 2.2, we deduce that
$\left\|u^{\rm E}_{t}-u_{0}\right\|_{B^{s-1}_{p,\infty}}\leq$ $~{}\int_{0}^{t}\left\|\mathcal{P}\left(u^{\rm E}_{\tau}\cdot\nabla u^{\rm E}_{\tau}\right)\right\|_{B^{s-1}_{p,\infty}}\mathrm{d}\tau$
$\leq$ $~{}\int_{0}^{t}\left\|\mathcal{P}\left(u^{\rm E}_{\tau}\cdot\nabla u^{\rm E}_{\tau}\right)\right\|_{\dot{B}^{s-1}_{p,\infty}\cap\dot{B}^{0}_{p,1}}\mathrm{d}\tau$
$\leq$ $~{}C\int_{0}^{t}\left\|u^{\rm E}_{\tau}\otimes u^{\rm E}_{\tau}\right\|_{\dot{B}^{s}_{p,\infty}\cap\dot{B}^{1}_{p,1}}\mathrm{d}\tau$
$\leq$ $~{}C\int_{0}^{t}\left\|u^{\rm E}_{\tau}\otimes u^{\rm E}_{\tau}\right\|_{B^{s}_{p,\infty}}\mathrm{d}\tau$
$\leq$ $~{}Ct\|u_{0}\|_{B^{s}_{p,\infty}}^{2}$
$\leq$ $~{}Ct\delta^{2},$ (3.7)
where we have used the boundedness of Riesz transforms in homogeneous Besov spaces and the fact that $B^{s}_{p,\infty}$ is a Banach algebra.
From (3.6), it follows that
$u^{\rm E}_{t}-u_{0}+t\mathcal{P}\left(u_{0}\cdot\nabla u_{0}\right)=\int_{0}^{t}\mathcal{P}\left(u_{0}\cdot\nabla u_{0}-u^{\rm E}_{\tau}\cdot\nabla u^{\rm E}_{\tau}\right)\mathrm{d}\tau.$
Thus, we obtain from (3.3) that
$\left\|u^{\rm E}_{t}-u_{0}+t\mathcal{P}\left(u_{0}\cdot\nabla u_{0}\right)\right\|_{\dot{B}^{s-2}_{p,\infty}}\leq$ $~{}C\int_{0}^{t}\left\|u_{0}\cdot\nabla u_{0}-u^{\rm E}_{\tau}\cdot\nabla u^{\rm E}_{\tau}\right\|_{B^{s-2}_{p,\infty}}\mathrm{d}\tau$
$\leq$ $~{}C\int_{0}^{t}\left\|u_{0}\otimes u_{0}-u^{\rm E}_{\tau}\otimes u^{\rm E}_{\tau}\right\|_{B^{s-1}_{p,\infty}}\mathrm{d}\tau$
$\leq$ $~{}C\int_{0}^{t}\left\|u^{\rm E}_{\tau}-u_{0}\right\|_{B^{s}_{p,\infty}}\|u_{0}\|_{B^{s}_{p,\infty}}\mathrm{d}\tau$
$\leq$ $~{}Ct^{2}\delta^{3}.$
Then we complete the proof of Proposition 3.3. ∎
Proposition 3.4.
Let $u_{0}$ be given by Lemma 3.1. Assume that $(s,p)$ satisfies (1.1). For $t\in(0,1)$, we have for some positive constant $C$ independent of $t,\varepsilon$ and $\delta$
$\left\|u^{\rm NS}_{\varepsilon}(t,u_{0})-e^{t\varepsilon\Delta}u_{0}+\int_{0}^{t}e^{(t-\tau)\varepsilon\Delta}\mathcal{P}\left(e^{\tau\varepsilon\Delta}u_{0}\cdot\nabla e^{\tau\varepsilon\Delta}u_{0}\right)\mathrm{d}\tau\right\|_{\dot{B}^{s-2}_{p,\infty}}\leq Ct^{2}\delta^{3}.$
Proof.
From now on, we set $u^{\rm NS}_{t}=u^{\rm NS}(t,u_{0})$ for simplicity. Due to the fact $B^{s}_{p,\infty}\hookrightarrow\rm Lip$, we know that there exists a positive time $T=T(\|u_{0}\|_{B^{s}_{p,\infty}})$ such that
$\|u^{\rm NS}_{t}\|_{L^{\infty}_{T}B^{s}_{p,\infty}}\leq C\|u_{0}\|_{B^{s}_{p,\infty}}\leq C\delta.$
By Duhamel’s principle, from (NS), one has
$u^{\rm NS}_{t}=e^{t\varepsilon\Delta}u_{0}-\int_{0}^{t}e^{(t-\tau)\varepsilon\Delta}\mathcal{P}\left(u^{\rm NS}_{\tau}\cdot\nabla u^{\rm NS}_{\tau}\right)\mathrm{d}\tau,$ (3.8)
which gives that
$\left\|u^{\rm NS}_{t}-e^{t\varepsilon\Delta}u_{0}\right\|_{B^{s-1}_{p,\infty}}\leq$ $~{}C\int_{0}^{t}\left\|u^{\rm NS}_{\tau}\cdot\nabla u^{\rm NS}_{\tau}\right\|_{B^{s-1}_{p,\infty}}\mathrm{d}\tau$
$\leq$ $~{}C\int_{0}^{t}\left\|u^{\rm NS}_{\tau}\otimes u^{\rm NS}_{\tau}\right\|_{B^{s}_{p,\infty}}\mathrm{d}\tau$
$\leq$ $~{}Ct\delta^{2}.$ (3.9)
From (3.8), we have
$u^{\rm NS}_{t}$ $-e^{t\varepsilon\Delta}u_{0}+\int_{0}^{t}e^{(t-\tau)\varepsilon\Delta}\mathcal{P}\left(e^{\tau\varepsilon\Delta}u_{0}\cdot\nabla e^{\tau\varepsilon\Delta}u_{0}\right)\mathrm{d}\tau$
$=-\int_{0}^{t}e^{(t-\tau)\varepsilon\Delta}\mathcal{P}\left(u^{\rm NS}_{\tau}\cdot\nabla u^{\rm NS}_{\tau}-e^{\tau\varepsilon\Delta}u_{0}\cdot\nabla e^{\tau\varepsilon\Delta}u_{0}\right)\mathrm{d}\tau,$
from which and Minkowski’s inequality, and using (3.3), we obtain that
$\left\|u^{\rm NS}_{t}-e^{t\varepsilon\Delta}u_{0}+\int_{0}^{t}e^{(t-\tau)\varepsilon\Delta}\mathcal{P}\left(e^{\tau\varepsilon\Delta}u_{0}\cdot\nabla e^{\tau\varepsilon\Delta}u_{0}\right)\mathrm{d}\tau\right\|_{\dot{B}^{s-2}_{p,\infty}}$
$\leq$ $~{}C\int_{0}^{t}\left\|\mathcal{P}\left(u^{\rm NS}_{\tau}\cdot\nabla u^{\rm NS}_{\tau}-e^{\tau\varepsilon\Delta}u_{0}\cdot\nabla e^{\tau\varepsilon\Delta}u_{0}\right)\right\|_{\dot{B}^{s-2}_{p,\infty}}\mathrm{d}\tau$
$\leq$ $~{}C\int_{0}^{t}\left\|u^{\rm NS}_{\tau}-e^{\tau\varepsilon\Delta}u_{0}\right\|_{B^{s-1}_{p,\infty}}\|u^{\rm NS}_{\tau}\|_{B^{s}_{p,\infty}}\mathrm{d}\tau$
$\leq$ $~{}Ct^{2}\delta^{3}.$
Then we complete the proof of Proposition 3.4. ∎
3.4 Non-Convergence
In this subsection, we shall prove the non-convergence of the solutions for the Navier-Stokes equations toward the Euler equations in the inviscid limit under the framework of weak Besov spaces $B^{s}_{p,\infty}$.
From (1.2) and (1.3), it follows that
$u^{\rm NS}_{t}-u^{\rm E}_{t}$ $=\left(e^{t\varepsilon\Delta}u_{0}-u_{0}\right)-\int_{0}^{t}e^{(t-\tau)\varepsilon\Delta}\mathcal{P}\left(e^{\tau\varepsilon\Delta}u_{0}\cdot\nabla e^{\tau\varepsilon\Delta}u_{0}-u_{0}\cdot\nabla u_{0}\right)\mathrm{d}\tau$
$\quad-\int_{0}^{t}\left(e^{\tau\varepsilon\Delta}-\mathrm{Id}\right)\mathcal{P}\left(u_{0}\cdot\nabla u_{0}\right)\mathrm{d}\tau-\left[u^{\rm E}_{t}-u_{0}+t\mathcal{P}\left(u_{0}\cdot\nabla u_{0}\right)\right]$
$\quad+u^{\rm NS}_{t}-e^{t\varepsilon\Delta}u_{0}+\int_{0}^{t}e^{(t-\tau)\varepsilon\Delta}\mathcal{P}\left(e^{\tau\varepsilon\Delta}u_{0}\cdot\nabla e^{\tau\varepsilon\Delta}u_{0}\right)\mathrm{d}\tau$
$=\left(e^{t\varepsilon\Delta}u_{0}-u_{0}\right)+\mathbf{I}_{1}+\mathbf{I}_{2}+\mathbf{I}_{3}+\mathbf{I}_{4},$ (3.10)
where
$\mathbf{I}_{1}:=-\int_{0}^{t}e^{(t-\tau)\varepsilon\Delta}\mathcal{P}\left(e^{\tau\varepsilon\Delta}u_{0}\cdot\nabla e^{\tau\varepsilon\Delta}u_{0}-u_{0}\cdot\nabla u_{0}\right)\mathrm{d}\tau,$
$\mathbf{I}_{2}:=-\int_{0}^{t}\left(e^{\tau\varepsilon\Delta}-\mathrm{Id}\right)\mathcal{P}\left(u_{0}\cdot\nabla u_{0}\right)\mathrm{d}\tau,$
$\mathbf{I}_{3}:=-\left[u^{\rm NS}_{t}-u_{0}+t\mathcal{P}\left(u_{0}\cdot\nabla u_{0}\right)\right],$
$\mathbf{I}_{4}:=u^{\rm NS}_{t}-e^{t\varepsilon\Delta}u_{0}+\int_{0}^{t}e^{(t-\tau)\varepsilon\Delta}\mathcal{P}\left(e^{\tau\varepsilon\Delta}u_{0}\cdot\nabla e^{\tau\varepsilon\Delta}u_{0}\right)\mathrm{d}\tau.$
Using the triangle inequality and Proposition 3.2, we have
$\left\|u^{\rm NS}_{t}-u^{\rm E}_{t}\right\|_{{B}^{s}_{p,\infty}}$ $\geq 2^{ns}\left\|{\Delta}_{n}\left(e^{t\varepsilon\Delta}u_{0}-u_{0}\right)\right\|_{L^{p}}-2^{ns}\sum_{i=1}^{4}\left\|{\Delta}_{n}\mathbf{I}_{i}\right\|_{L^{p}}$
$\geq\delta 2^{n}t(c-Ct2^{n})-C2^{n}\left(\left\|\mathbf{I}_{1}\right\|_{\dot{B}^{s-1}_{p,\infty}}+\left\|\mathbf{I}_{2}\right\|_{\dot{B}^{s-1}_{p,\infty}}\right)$
$\quad-C2^{2n}\left(\left\|\mathbf{I}_{3}\right\|_{\dot{B}^{s-2}_{p,\infty}}+\left\|\mathbf{I}_{4}\right\|_{\dot{B}^{s-2}_{p,\infty}}\right).$ (3.11)
Using Proposition 3.2, we have
$\left\|\mathbf{I}_{1}\right\|_{\dot{B}^{s-1}_{p,\infty}}+\left\|\mathbf{I}_{2}\right\|_{\dot{B}^{s-1}_{p,\infty}}\leq Ct\delta^{2}.$
Using Propositions 3.3-3.4, we have
$\left\|\mathbf{I}_{3}\right\|_{\dot{B}^{s-2}_{p,\infty}}+\left\|\mathbf{I}_{4}\right\|_{\dot{B}^{s-2}_{p,\infty}}\leq Ct^{2}\delta^{3}.$
Inserting the above into (3.4), and picking $t=\delta 2^{-n}$ with the above $\delta$, we deduce that for large $n$
$\left\|u^{\rm NS}_{t}-u^{\rm E}_{t}\right\|_{{B}^{s}_{p,\infty}}$ $\geq\delta 2^{n}t\left(c-Ct2^{n}\right)-C\delta^{2}t2^{n}-C\delta^{3}t^{2}2^{2n}$
$\geq\delta^{2}\left(c-C\delta\right)-C\delta^{3}-C\delta^{5}.$
Thus, choosing $\delta$ sufficiently small, we deduce that
$\left\|u^{\rm NS}_{\varepsilon}(t,u_{0})-u^{\rm E}(t,u_{0})\right\|_{B^{s}_{p,\infty}}$ $\geq c_{0}\delta^{2}.$
Notice that $t\to 0^{+}$ and $\varepsilon\to 0^{+}$ as $n\to\infty$, we complete the proof of Theorem 1.2. $\square$
Declarations
Availability of data and materials
No data was used for the research described in the article.
Conflict of interest
The authors declare that they have no conflict of interest.
Funding
J. Li is supported by the National Natural Science Foundation of China (No.12161004), Innovative High end Talent Project in Ganpo Talent Program (No.gpyc20240069), Training Program for Academic and Technical Leaders of Major Disciplines in Ganpo Juncai Support Program (No.20232BCJ23009).